EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Binomial & Poisson Distributions · Question 25 of 68

Question 25

MediumOriginal question

A medicine cures 70% of patients. It is given to 3 patients. What is the probability that exactly 2 are cured?

Options for question 25

Show the answer and explanation
CorrectNot quiteAnswer D

Correct answer: D

Explanation

  1. P(X=2)=3C2(0.7)2(0.3)P(X = 2) = {}^{3}C_2(0.7)^{2}(0.3).

  2. =3×0.49×0.3=0.441= 3 \times 0.49 \times 0.3 = 0.441.

Worth noting

0.147 forgets the 3 ways to choose which two patients are cured.

Concept
Exactly r successes
Typical time
90 s
Your time
—
The rule this testsBinomial probabilities
Practise this chapter →

Ten questions from Binomial & Poisson Distributions, at your own pace.

Practise this chapter →

Ten questions from Binomial & Poisson Distributions, at your own pace.